# Blade: A Derivative-free Bayesian Inversion Method using Diffusion Priors

### v1

@@ B.2 Prior step
The prior step is implemented as a denoising diffusion process, with its pseudocode detailed in Algorithm 3. We set $\sigma(t)=t$ for simplicity and employ the Euler ODE sampler for faster sampling. We discretize the denoising diffusion process with the standard Euler method. Following Karras et al. (2022), we use the following step size:
@@ B.3 Annealing schedule
Inspired by the discretization scheme proposed by Karras et al. (2022), we consider the following schedule, which we refer to as EDM schedule:
@@ C.2.1 Problem setup
Following the experimental setup in InverseBench (Zheng et al., 2025b), we consider the 2-d Navier-Stokes equation for a viscous, incompressible fluid in vorticity form on a torus,
@@
where $\bm{u}$} is the velocity field, ${\bm{w}}=\nabla\times{\bm{u}}$ is the vorticity, ${\bm{w}}_{0}$ is the initial vorticity, $\nu\in\mathbb{R}_{+}$ is the viscosity coefficient, and $f$ is the forcing function. The solution operator $\mathcal{F}$ is defined as the operator mapping the vorticity from the initial vorticity to the vorticity at time $T$. $\mathcal{F}:{\bm{w}}_{0}\rightarrow{\bm{w}}_{T}$. Numerically, it is realized as a pseudo-spectral solver (He & Sun, 2007). This Navier-Stokes equation is a standard benchmark problem widely used in the literature (Iglesias et al., 2013; Li et al., 2020; Takamoto et al., 2022). The forward model in our inverse problem is given by
@@ C.2.2 Evaluation metrics
The CRPS (Gneiting & Raftery, 2007) is a standard probabilistic metric to assess the quality of the entire predicted distribution for inverse problems, which is defined as
@@
where ${\bm{x}},{\bm{x}}^{\prime}$ are independent random predictions and $\bm{x}$$^{*}$ is the single observed ground truth. Intuitively, it measures the distance between a predicted distribution and the single observed ground truth $\bm{x}$$^{*}$ that actually occurred. It is minimized when the ensemble prediction is drawn from the same distribution as the ground truth, i.e., ${\bm{x}}^{(j)}\sim p({\bm{x}}^{*})$ for all $j$. We consider the multi-dimensional version of CRPS defined in Rasp et al. (2024). For an ensemble prediction $\{{\bm{x}}^{(j)}\}_{j=1}^{J}$ where ${\bm{x}}^{(j)}\in\mathbb{R}^{n}$, the CRPS for the single ground truth $\bm{x}$$^{*}$ is given by
@@
which can be implemented in $O(nJ\log J)$ complexity using the equivalent form introduced in Zamo & Naveau (2018). In our experiments, we report the CRPS averaged over all test cases.
@@
The spread-skill ratio (SSR) is a simple yet powerful diagnostic of how well an ensemble prediction’s stated uncertainty (spread) matches its actual error (skill) (Fortin et al., 2014). Intuitively, if the ensemble distribution truly captures the variability of the ground truth, then ensemble members should be statistically indistinguishable from observed outcomes. Formally, let $\{{\bm{x}}^{*}_{i}\}_{i=1}^{N}$ denote a set of observed ground truths. Suppose, for each observed ground truth ${\bm{x}}^{*}_{i}$, we have an ensemble prediction $\{{\bm{x}}_{i,j}\}_{j=1}^{J}$. Let $\bar{{\bm{x}}}_{i}=\frac{1}{J}\sum_{j}{\bm{x}}_{i,j}$. The unbiased estimator of SSR can be written as
@@
The rank histogram (Anderson, 1996; Hamill, 1997; Talagrand, 1999) assesses ensemble calibration by comparing the truth to the empirical distribution formed by the ensemble. For each grid point in each test case, the ensemble members are sorted, and the rank of the true value within this ordering (0 means below all, $J$ means above all, or an intermediate integer) is recorded. Pooling these ranks over all points and cases yields a histogram with $J+1$ bins. A flat histogram indicates a statistically consistent ensemble: the truth behaves like an additional random draw from the predictive distribution. A U-shape signals under-dispersion (ensemble spread too narrow), an inverted U indicates over-dispersion, and tilted shapes reveal bias.
@@ C.2.3 Baseline implementation
For baseline methods that do not require additional training on paired data, such as DPG (Tang et al., 2024), SCG (Huang et al., 2024), EnKG (Zheng et al., 2025a), EKI (Iglesias et al., 2013), EKS (Garbuno-Inigo et al., 2020a) (with diffusion prior initialization), we follow the implementation provided in InverseBench (Zheng et al., 2025b). For methods that do require training on paired data, specifically the end-to-end UNet and conditional diffusion model (CDM), we first generate a collection of observation-solution pairs by simulating observations from the prior training dataset available in InverseBench (Zheng et al., 2025b). To evaluate their in-distribution performance, we retrain the UNet and CDM for each noise level, which takes around 7-10 hours on a single GH200.
The end-to-end UNet architecture was adapted from the UNet used in our diffusion model by removing the time conditioning branch. The observation is upsampled to the same resolution before being fed into the UNet. However, it is important to note that observations are not always spatially aligned with the unknown signal in a general setting. Consequently, end-to-end neural networks typically require additional design considerations for different types of observations.
The CDM was also adapted from the UNet architecture of the prior diffusion model. This involved replacing the self-attention module with cross-attention and incorporating a CNN-based observation encoder, following the conditioning mechanism used in Rombach et al. (2022).
@@ C.3.1 Problem setup
We evaluated our method on image restoration tasks with the FFHQ256 dataset. Our evaluation set consisted of the first ten images, indexed 00000 to 00009, in the validation set. The pre-trained model is taken from Chung et al. (2023) (FFHQ256) and converted into an EDM checkpoint with their Variance-Preserving (VP) preconditioning (Karras et al., 2022). In general, we follow the experiment setup of  Chung et al. (2023). All problem settings use a measurement noise of $\sigma_{y}=0.05$. We address that PnP-DM results on phase retrieval are significantly worse than those originally reported. Note that we used a different forward model configuration, larger measurement noise, and full-color images, which differs from the original PnP-DM setup. Furthermore, PnP-DM super-resolution results are lower than originally reported. We note that we compare against the PnP-DM configuration that uses Langevin Monte Carlo during the likelihood step, which differs from their original configuration for linear inverse problems.
@@ C.3.2 Evaluation metrics
The structural similarity index measure (SSIM)  (Wang et al., 2004) is another metric to compute the similarity between two images. It compares patterns of luminance, contrast, and structure between two images to achieve a metric more aligned with visual perception, given by
@@
The Learned Perceptual Image Patch Similarity (LPIPS)  (Zhang et al., 2018) compares two images by measuring their distance in the feature space of a pre-trained neural network. LPIPS is computed as a weighted sum of the Euclidean distances between the network activations of an image $\bm{x}$} and an image $\bm{x}$$^{*}$.
@@ Appendix D Additional experiments
We provide additional quantitative results on linear Gaussian problems. We report the SWD and $D_{\mathrm{KL}}$ for each method across different problem dimensions and observation noise levels in Table 4. As shown, Blade, EKS and MCGDiff (Cardoso et al., 2024) both achieve similar and strong performance, surpassing the other tested baselines. This aligns with theoretical expectations, given that these methods are proven to be asymptotically accurate in linear Gaussian problems. Consequently, the results empirically confirm that Blade can achieve accurate posterior sampling when applied to linear problems.
@@
We also represent results of Blade on black hole imaging problem from InverseBench (Zheng et al., 2025b) in Table 3. As shown, Blade attains the best reconstruction quality (PSNR) among compared methods while remaining competitive on the EHT closure statistics. Additionally, we note that InverseBench assumes a single ground truth and does not evaluate posterior calibration or multi-modality. Since the goal in this work is calibrated posterior sampling, we view these additional results as complementary evidence.
Table 3: Evaluating Blade on the black hole imaging benchmark from InverseBench (Zheng et al., 2025b). Note that this benchmark only uses deterministic metrics with a focus on accurate point estimate. Blade’s goal is well-calibrated posterior sampling. We therefore regard the results as complementary evidence.
@@
Figure 10: Scaling of different methods’ performance with training data size across three measurement noise levels. Each row corresponds to one evaluation metric. UNet produces deterministic prediction so the probabilistic metrics like CRPS and SSR do not apply. Note that our method uses the same diffusion prior across different noise levels while the training-based methods train separate networks for each noise level.

### v2

@@ B.2 Prior step
The prior step is implemented as a denoising diffusion process, with its pseudocode detailed in Algorithm 3. We set $\sigma(t)=t$ for simplicity and employ the Euler ODE sampler for faster sampling. We discretize the denoising diffusion process with the standard Euler method. Following karras2022elucidating, we use the following step size:
@@ B.3 Annealing schedule
Inspired by the discretization scheme proposed by karras2022elucidating, we consider the following schedule, which we refer to as EDM schedule:
@@ C.2.1 Problem setup
Following the experimental setup in InverseBench (zhenginversebench), we consider the 2-d Navier-Stokes equation for a viscous, incompressible fluid in vorticity form on a torus,
@@
where $\bm{u}$} is the velocity field, ${\bm{w}}=\nabla\times{\bm{u}}$ is the vorticity, ${\bm{w}}_{0}$ is the initial vorticity, $\nu\in\mathbb{R}_{+}$ is the viscosity coefficient, and $f$ is the forcing function. The solution operator $\mathcal{F}$ is defined as the operator mapping the vorticity from the initial vorticity to the vorticity at time $T$. $\mathcal{F}:{\bm{w}}_{0}\rightarrow{\bm{w}}_{T}$. Numerically, it is realized as a pseudo-spectral solver (he2007stability). This Navier-Stokes equation is a standard benchmark problem widely used in the literature (iglesias2013ensemble; li2020fourier; takamoto2022pdebench). The forward model in our inverse problem is given by
@@ C.2.2 Evaluation metrics
The CRPS (gneiting2007strictly) is a standard probabilistic metric to assess the quality of the entire predicted distribution for inverse problems, which is defined as
@@
where ${\bm{x}},{\bm{x}}^{\prime}$ are independent random predictions and $\bm{x}$$^{*}$ is the single observed ground truth. Intuitively, it measures the distance between a predicted distribution and the single observed ground truth $\bm{x}$$^{*}$ that actually occurred. It is minimized when the ensemble prediction is drawn from the same distribution as the ground truth, i.e., ${\bm{x}}^{(j)}\sim p({\bm{x}}^{*})$ for all $j$. We consider the multi-dimensional version of CRPS defined in rasp2024weatherbench. For an ensemble prediction $\{{\bm{x}}^{(j)}\}_{j=1}^{J}$ where ${\bm{x}}^{(j)}\in\mathbb{R}^{n}$, the CRPS for the single ground truth $\bm{x}$$^{*}$ is given by
@@
which can be implemented in $O(nJ\log J)$ complexity using the equivalent form introduced in zamo2018estimation. In our experiments, we report the CRPS averaged over all test cases.
@@
The spread-skill ratio (SSR) is a simple yet powerful diagnostic of how well an ensemble prediction’s stated uncertainty (spread) matches its actual error (skill) (fortin2014should). Intuitively, if the ensemble distribution truly captures the variability of the ground truth, then ensemble members should be statistically indistinguishable from observed outcomes. Formally, let $\{{\bm{x}}^{*}_{i}\}_{i=1}^{N}$ denote a set of observed ground truths. Suppose, for each observed ground truth ${\bm{x}}^{*}_{i}$, we have an ensemble prediction $\{{\bm{x}}_{i,j}\}_{j=1}^{J}$. Let $\bar{{\bm{x}}}_{i}=\frac{1}{J}\sum_{j}{\bm{x}}_{i,j}$. The unbiased estimator of SSR can be written as
@@
The rank histogram (anderson1996method; hamill1997reliability; talagrand1999evaluation) assesses ensemble calibration by comparing the truth to the empirical distribution formed by the ensemble. For each grid point in each test case, the ensemble members are sorted, and the rank of the true value within this ordering (0 means below all, $J$ means above all, or an intermediate integer) is recorded. Pooling these ranks over all points and cases yields a histogram with $J+1$ bins. A flat histogram indicates a statistically consistent ensemble: the truth behaves like an additional random draw from the predictive distribution. A U-shape signals under-dispersion (ensemble spread too narrow), an inverted U indicates over-dispersion, and tilted shapes reveal bias.
@@ C.2.3 Baseline implementation
For methods that require training on paired data, specifically the end-to-end U-Net and conditional diffusion model (CDM), we first generate a collection of observation-solution pairs by simulating observations from the prior training dataset available in InverseBench (zhenginversebench). To evaluate their in-distribution performance, we retrain the U-Net and CDM for each noise level, which takes around 7-10 hours on a single GH200.
U-Net
The end-to-end U-Net baseline is adapted from the U-Net used in our diffusion model by removing the time conditioning branch. The observation is upsampled to the same resolution before being fed into the U-Net. However, it is important to note that observations are not always spatially aligned with the unknown signal in a general setting. Consequently, end-to-end neural networks typically require additional design considerations for different types of observations.
CDM-CA
CDM-CA is adapted from the U-Net architecture of the prior diffusion model. This involved replacing the self-attention module with cross-attention and incorporating a CNN-based observation encoder, following the conditioning mechanism used in rombach2022high.
CDM-Cat
CDM-Cat is adapted from the U-Net of the prior diffusion model. We upsample the observation to the same resolution as the solution and directly concatenate it with the input over the channel dimension. This conditioning mechanism only works for observations that are spatially aligned to the input.
To find the right hyperparameters for the training-free methods, we perform two rounds of Bayesian optimization over their hyperparameters on the validation set provided in InverseBench (zhenginversebench).
Existing baselines in InverseBench
We simply take the existing implementations from InverseBench (zhenginversebench) for baseline methods that already exist in the codebase including DPG (tang2024solving), SCG (huang2024symbolic), EnKG (zheng2024ensemble), and EKI (iglesias2013ensemble).
EKS + DM
The original EKS (garbuno2020interacting) method only considers the Gaussian prior. The most straightforward way to incorporate the diffusion prior is to initialize the ensemble members with random samples from diffusion model. We fix the ensemble size to 1024 and optimize the other hyperparameters using Bayesian optimization.
EKS w/ DP
While the original update rules of EKS is derived for Gaussian prior only, we can approximate the gradient of the prior regularization term in EKS by evaluating the diffusion model $s_{\theta}({\bm{x}}_{t};\sigma(t))$ with $t$ close to zero. We fix the ensemble size to 1024 and optimize the other hyperparameters using Bayesian optimization.
Localized EKS w/ DP
Localization is a common technique to improve the multi-modal sampling of ensemble methods by assigning a localized mean and covariance for each particle. We implement the localized EKS w/ DP following the formulation in  (reich2021fokker; wagner2022ensemble). Again, we fix the ensemble size to 1024 and optimize the other hyperparameters using Bayesian optimization.
FKD
Feynman-Kac diffusion steering (FKD) (zhao2025conditional; singhal2025general) is a method based on interacting particle system, which can be viewed as a generalization of SCG (huang2024symbolic). We fix the ensemble size to 1024 and optimize the other hyperparameters using Bayesian optimization.
@@ C.3.1 Problem setup
We evaluated our method on image restoration tasks with the FFHQ256 dataset. Our evaluation set consisted of the first ten images, indexed 00000 to 00009, in the validation set. The pre-trained model is taken from chung2023diffusion (FFHQ256) and converted into an EDM checkpoint with their Variance-Preserving (VP) preconditioning (karras2022elucidating). In general, we follow the experiment setup of  chung2023diffusion. All problem settings use a measurement noise of $\sigma_{y}=0.05$. We address that PnP-DM results on phase retrieval are significantly worse than those originally reported. Note that we used a different forward model configuration, larger measurement noise, and full-color images, which differs from the original PnP-DM setup. Furthermore, PnP-DM super-resolution results are lower than originally reported. We note that we compare against the PnP-DM configuration that uses Langevin Monte Carlo during the likelihood step, which differs from their original configuration for linear inverse problems.
@@ C.3.2 Evaluation metrics
The structural similarity index measure (SSIM)  (wang2004image) is another metric to compute the similarity between two images. It compares patterns of luminance, contrast, and structure between two images to achieve a metric more aligned with visual perception, given by
@@
The Learned Perceptual Image Patch Similarity (LPIPS)  (zhang2018unreasonable) compares two images by measuring their distance in the feature space of a pre-trained neural network. LPIPS is computed as a weighted sum of the Euclidean distances between the network activations of an image $\bm{x}$} and an image $\bm{x}$$^{*}$.
@@ Appendix D Additional experiments
We provide additional quantitative results on linear Gaussian problems. We report the SWD and $D_{\mathrm{KL}}$ for each method across different problem dimensions and observation noise levels in Table 4. As shown, Blade, EKS and MCGDiff (cardoso2024monte) both achieve similar and strong performance, surpassing the other tested baselines. This aligns with theoretical expectations, given that these methods are proven to be asymptotically accurate in linear Gaussian problems. Consequently, the results empirically confirm that Blade can achieve accurate posterior sampling when applied to linear problems.
@@
We also represent results of Blade on black hole imaging problem from InverseBench (zhenginversebench) in Table 3. As shown, Blade attains the best reconstruction quality (PSNR) among compared methods while remaining competitive on the EHT closure statistics. Additionally, we note that InverseBench assumes a single ground truth and does not evaluate posterior calibration or multi-modality. Since the goal in this work is calibrated posterior sampling, we view these additional results as complementary evidence.
Table 3: Evaluating Blade on the black hole imaging benchmark from InverseBench (zhenginversebench). Note that this benchmark only uses deterministic metrics with a focus on accurate point estimate. Blade’s goal is well-calibrated posterior sampling. We therefore regard the results as complementary evidence.
@@
Figure 10: Scaling of different methods’ performance with training data size across three measurement noise levels. Each row corresponds to one evaluation metric. U-Net produces deterministic prediction so the probabilistic metrics like CRPS and SSR do not apply. Note that our method uses the same diffusion prior across different noise levels while the training-based methods train separate networks for each noise level.